Rachel is 'x' years old. Sophie is double Rachel's age. Tony is 3 years younger than Rachel. If the total of their ages is 41. How old is Rachel?
Answer by ankor@dixie-net.com(22740)

Rachel is 'x' years old.
r = x
Sophie is double Rachel's age.
s = 2x
Tony is 3 years younger than Rachel.
t = (r-3)
:
If the total of their ages is 41.
r + s + t = 41
How old is Rachel?
x + 2x + (x-3) = 41
4x = 41 + 3
x - 44/4
x = 11 yr is Rachel's age
:
:
See if that checks out, s = 22, t = 8
11 + 22 + 8 = 41